An aeroplane moves horizantaly at 162km/h speed at height of 2km above the ground, towards a fixed Target on a ground. If a bomb is to released to hit the target from the aeroplane detarmine where should the bomb be released? Also calculate the velocity with which the bomb strike the target
1
Expert's answer
2019-10-14T09:31:47-0400

To cover the vertical distance, "s=u\times t+\frac{1}{2}a\times t^2"

"s=2Km=2000m"

"2=\frac{1}{2}\times9.8\times t^2"

"t=20.20s"

So horizontal distance at which the bomb should be dropped"=v\times t=45\times20.20=909m"

Speed at which bomb is dropped"=9.8\times t=9.8\times20.20=197.96\frac{m}{sec}"